OpenStudy (anonymous):

θ=2 find measure of θin terms of degrees, minutes, seconds.

OpenStudy (anonymous):

\[\theta=2\pi?\]

OpenStudy (jdoe0001):

\(\bf degrees=\cfrac{{\color{brown}{ radians}}\cdot 180}{\pi}\to \cfrac{{\color{brown}{ 2}}\cdot 180}{\pi}\approx 114.{\color{blue}{ 59}}^o \\ \quad \\ \textit{there are 60 minutes in 1 degree, how many in 0.59?} \\ \quad \\ \begin{array}{ccllll} minute&degree \\\hline\\ 60&1\\ x&0.59 \end{array}\to \cfrac{60}{x}=\cfrac{1}{{\color{blue}{ 0.59}}}\)

OpenStudy (jdoe0001):

and whatever you get in minutes, will be \(\bf 144^o\square'\) if the amount gives you a decimal, then you'd need to get the seconds to get the the seconds, you'd do the same \(\large \begin{array}{ccllll} seconds&minutes \\\hline\\ 60&1\\ x&\textit{decimal amount here} \end{array}\)

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

yw